A 2.00 L sample of gas originally at 25 °C and pressure of 400 torr is allowed to expand
to a volume of 5.00 L.
If the final pressure of gas is 585 torr, what is the final temperature in K?
1
Expert's answer
2020-12-06T13:28:47-0500

p*V = n*R*T

n and R are constant so p*V/T = constant

Then (2.00 L)*(400 torr)/(25+273.15 K) = (5.00 L)*(585 torr)/T

T = (5.00 L)*(585 torr)*(25+273.15 K)/(2.00 L)/(400 torr) = 1090 K